OpenStudy (anonymous):

why is the product of a negative integer and a positive integer a negative integer? I know basic math skills and I know that (-)x(+)=(-) but how do I explain to someone why (-)x(+)=(-)?

OpenStudy (anonymous):

depends on what you get to assume

OpenStudy (anonymous):

like if you were dealing with someone who knew nothing about math whatsoever, how would you teach them and get them to understand that a (-)x(+)=(-) and (-)x(-)=(+)?

OpenStudy (anonymous):

the connection between plus and minus (which live in the world of addition) and multiplication always comes from the distributive property

OpenStudy (anonymous):

okay so I was able to find a way to explain that, but how would you explain that the product of two negative numbers is positive?

OpenStudy (anonymous):

in particular, how would you convince a skeptic that -5x-6=30?

OpenStudy (anonymous):

you could look at \[5\times 6+(-5)\times 6+(-5)\times (-6)\] and prove that \(5\times 6)=(-5)\times (-6)\] it might be more trouble than it is worth, but it is not hard

OpenStudy (anonymous):

\[5\times 6+(-5)\times 6+(-5)\times -6)\] can be rewritten as \[5\times 6+(-5)\times (6+(-6)]\] by the distributive law, and since \(6+(-6)=0\) this is the same as \[5\times 6\] at the same time it is \[5\times (6+(-6)+(-5)\times (-6)=0+(-5)\times (-6)=(-5)\times (-6)\]

OpenStudy (anonymous):

i seem to have missed some parentheses there in any case you get that it is both \(5\times 6\) and \(-5)\times (-6)\) so they are equal`

geerky42 (geerky42):

Think it as walking forward or backward and facing positive or negative direction. When you just walk forward at positive direction, you just go forward, which is just \((+)(+)=+\) When you walk backward and facing positive direction, you basically go backward \((-)(+)=-\) When you walk forward facing negative direction, you also go backward. \((+)(-)=-\) BUT when you walk backward facing negative direction, you end up moving in positive direction. That's \((-)(-)=+\)

geerky42 (geerky42):

Just explanation in layman's terms.

OpenStudy (anonymous):

thank you all so much! so in terms of (-)x(-)=(+), i have this problem that is confusing me... Look for a pattern: Multiply 3x-6, 2x-6, 1x-6, 0x-6. Based on the pattern you see, what should -1x-6 be? What about -2x-6? -3x-6? -5x-6?

geerky42 (geerky42):

Start with easier pattern. \(~~~3\times-1=-3\) \(~~~2\times-1=-2\) \(~~~1\times-1=-1\) \(~~~0\times-1=0\) \(-1\times-1=1\) \(-2\times-1=2\) \(-3\times-1=3\)

OpenStudy (anonymous):

oh duh! I completely overlooked that and was just overwhelmed with all the numbers!

geerky42 (geerky42):

Another explanation is when I say "Eat!" I am encouraging you to eat (positive) But when I say "Do not eat!" I am saying the opposite (negative). Now if I say "Do NOT not eat!", I am saying I don't want you to starve, so I am back to saying "Eat!" (positive).

OpenStudy (anonymous):

thank you so much! :) great explanation!!

geerky42 (geerky42):

Welcome!

geerky42 (geerky42):

We can understand that negative times negative is hard concept to grasp.

OpenStudy (perl):

*